OpenStudy (zyberg):

Is my solution right? Problem: Two cars start to constantly accelerate from one point to the same direction. After time t_1 the distance between them is l. How much time t_2 would need to pass, so that the distance between the cars would be equal to 3l?

OpenStudy (zyberg):

\(a = \frac{v}{t}\) \(v = \frac{s}{t}\) so, \(a = \frac{s}{t^2}; s = at^2\) Let's say that a_2 > a_1; Then we make a system of equations: \(a_2 t_1^2 - a_1 t_1^2 = l\) -> \(t_1^2(a_2 - a_1) = l\) \(a_2 t_2^2 - a_1 t_2^2 = 3l\) -> \(t_2^2(a_2 - a_1) = 3l\) then we just put the l from the first equation into the second and solve for t_2 to get: \(t_2 = \sqrt{3t_1^2} = t_1\sqrt{3}\) Is that the right way to solve this problem? @ganeshie8

OpenStudy (irishboy123):

\(\color{green}{\Huge \checkmark }\)

ganeshie8 (ganeshie8):

Your formula ` v = s/t ` doesn't look correct. Aren't you wrongly assuming that the velocity is constant ?

OpenStudy (zyberg):

@ganeshie8 you are right, however, how could I get distance from only acceleration and time?

ganeshie8 (ganeshie8):

We use the given info that acceleration is constant

ganeshie8 (ganeshie8):

How good are you with calculus - integration and differentiation ?

OpenStudy (zyberg):

Only basics.

ganeshie8 (ganeshie8):

Basics will do. What's the relation between acceleration and velocity ?

OpenStudy (zyberg):

a = v/t

ganeshie8 (ganeshie8):

That works only when acceleration is constant. Mayble let me ask this first : What's the relation between velocity and displacement ?

OpenStudy (zyberg):

Maybe dv/dt?

OpenStudy (zyberg):

Though that's exactly the same...

ganeshie8 (ganeshie8):

Yes, a = dv/dt

ganeshie8 (ganeshie8):

Next, what's the relation between velocity and displacement ?

OpenStudy (zyberg):

v = s/t

ganeshie8 (ganeshie8):

Again, that works only when velocity is constant. Think..

OpenStudy (zyberg):

v ds/dt

ganeshie8 (ganeshie8):

You mean v = ds/dt

OpenStudy (zyberg):

yes.

ganeshie8 (ganeshie8):

Imagine you're driving and your displacement from your home is given by the equation s = 3t^2 What is the function that gives your velocity at ANY point in time ?

OpenStudy (zyberg):

Well, it would be F(t) = 3t^2

ganeshie8 (ganeshie8):

s = 3t^2 is your displacement from the home. How can your velocity be same as displacement ?

OpenStudy (zyberg):

Ups. Sorry. I still haven't waken up. s/3t^2 = v

ganeshie8 (ganeshie8):

Haha I see you need some good coffee :) s = 3t^2 implies s/3t^2 = 1, not s/3t^2 = v

OpenStudy (zyberg):

Oh. Yes, I must have a cup. :D so, it would be the 3t^2/t = v; 3t = v

ganeshie8 (ganeshie8):

Wrong. As I've said earlier, there is no reason to assume that velocity is constant. v = s/t works only when velocity is constant.

OpenStudy (zyberg):

But how would we find the d of 3t^2?

ganeshie8 (ganeshie8):

Yes, we must use the relation between velocity and displacement : v = ds/dt

OpenStudy (zyberg):

(0 - 3t^2)/(0 -t) ?

ganeshie8 (ganeshie8):

For now memorize this formula : \[\large \dfrac{d}{dt} (t^n) = n*t^{n-1}\]

ganeshie8 (ganeshie8):

\[\large \dfrac{d}{dt} (t^2) = ?\]

OpenStudy (zyberg):

2t

ganeshie8 (ganeshie8):

\[\large \dfrac{d}{dt} (3t^2) = ?\]

OpenStudy (zyberg):

6t

ganeshie8 (ganeshie8):

good

ganeshie8 (ganeshie8):

So, if I give you a function for displacement, you can find the corresponding function for velocity ?

ganeshie8 (ganeshie8):

I give you `displacement` you can give me `velocity `

ganeshie8 (ganeshie8):

good so far. what if i gave you the velocity function and asked you to find the corresponding displacement function ?

OpenStudy (zyberg):

ds/dt = v ds = dt * v

ganeshie8 (ganeshie8):

I want "s", not "ds"

OpenStudy (zyberg):

I have no idea how to cancel out the d ;)

ganeshie8 (ganeshie8):

Take the previous function itself : v = 6t I want you to recover the corresponding displacement function. (assume your starting velocity is 0)

ganeshie8 (ganeshie8):

I know you can't do it yet.. let's take a quick look at a more simpler problem

OpenStudy (zyberg):

Alright :)

ganeshie8 (ganeshie8):

Suppose you're cycling at a constant velocity of 6 m/s. How much distance would you be travelling in 5 seconds ?

OpenStudy (zyberg):

v*s = 30 meters

ganeshie8 (ganeshie8):

Let's work it in another way (this method is going to make you love calculus)

ganeshie8 (ganeshie8):

Let's first plot the velocity function : v(t) = 6